OpenStudy (swag):

What is the sum of the interior angles of a nonagon? Answer 1260° 1440° 1620° 1800°

OpenStudy (experimentx):

1260

OpenStudy (callisto):

sum of interior angle = (n-2)*180 = (9-2)*180 = 7*180 = 1260 degree